OpenStudy (anonymous):

EValuate 5^-2 + 5 +5^0

OpenStudy (anonymous):

something to the negative power is the same as 1/5^2 know what I mean? and something to the power of 0 is 1 show me what you get and I will tell you if it is right.

OpenStudy (anonymous):

confused O.o

OpenStudy (mertsj):

\[5^{-2}+5+5^{0}=\frac{1}{5^{2}}+5+1=\frac{1}{25}+6=6 \frac{1}{25}\]

OpenStudy (anonymous):

5^-2=0.05, 5 and 5^0=1 so, its addition will be 6.05

OpenStudy (anonymous):

x^-1 is the same as 1/x^1 and x^-2 is the same as 1/x^2 understand? x^0 is 1 no matter what x is so 5^0 is 1. Use these facts to solve the problem.

OpenStudy (mertsj):

@Ravi2 5^-2 is NOT .05

OpenStudy (anonymous):

It's 1/25

OpenStudy (mertsj):

Yes. Which is .04

OpenStudy (anonymous):

ohhhhhhhhhhh!! yeah

OpenStudy (anonymous):

haha it's an easy mistake to make

OpenStudy (anonymous):

it made mistake because of continuously since 10 hours I sit @ computer../internet